OpenStudy (anonymous):

help me to solve this: A block of ice in the shape of a cube originally having volume 1000cm3 is melting in such way that the length of each of its edges is decreasing at the rate of 2cm/hr. At what rate is its surface area decreasing at the time its volume is 64cm3? Assume that the block of ice maintains its cubical shape? [ans: -96cm2/hr] show me the full solution.

OpenStudy (dumbcow):

Let s be length of an edge. \[V = s^{3}\] \[SA = 6s^{2}\] \[\frac{ds}{dt} = -2\] Find dA/dt, where dA refers to change in surface area \[\frac{dA}{dt} = \frac{ds}{dt}*\frac{dA}{ds}\] \[\frac{dA}{ds} = 12s\] \[\frac{dA}{dt} = (-2)(12s) = -24s\] We want the rate when V=64, find the s that gives a volume of 64 \[s^{3} = 64\rightarrow s = 4\] \[\frac{dA}{dt} = -24*4 = -96\]

OpenStudy (anonymous):

thanks..

OpenStudy (dumbcow):

no problem

OpenStudy (anonymous):

can i know, how about the volume 1000cm^3 given?